OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences.

OpenStudy (anonymous):

A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

@Loser66 i need help w. explaining the transformation

OpenStudy (loser66):

ok, you need understand how to graph those functions from the original one, right?

OpenStudy (anonymous):

no i need to understand the effect of the transformation on the last question @Loser66

OpenStudy (loser66):

I think to explain it, you need all in the same coordinate.

OpenStudy (anonymous):

??? @Loser66

OpenStudy (anonymous):

https://www.desmos.com/calculator/0zzwz6u4gk this is the graph

OpenStudy (anonymous):

now explain it to me please what is so confusing

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

ok, so you have D is \(f(x) = 23(1.06)^x\) right?

OpenStudy (loser66):

nope, hehehe. it is \(f(x) =23(1.06)^x -6\)

OpenStudy (loser66):

now, f(x+2) that is just x plus 2, hence \(f(x+2) =23(1.06)^{x+2}-6\)

OpenStudy (anonymous):

JUST EXPLAIN ME THE EFFECT OF THE TRASNFORMATIONS

OpenStudy (anonymous):

WHAT IS THE EFFFECT OF THIS OE F(X+2)

OpenStudy (loser66):

\(= 23(1.06)^x(1.06)^2 -6=25.8428(1.06)^x-6\)

OpenStudy (loser66):

calm down, no need to shout at me, hehehe. step by step

OpenStudy (loser66):

I graph them and you can see what effect when your graphs changing

OpenStudy (loser66):

https://www.desmos.com/calculator/4q6mdevevr

OpenStudy (loser66):

see, the orange graph is increasing faster than the red one when x =-50

OpenStudy (loser66):

now, it is f(-x) the green graph shows that the profit decreases as x goes to the right side. That is the profit is decreasing as time past https://www.desmos.com/calculator/bkg1zz7vvi

OpenStudy (loser66):

sorry, above is the blue one

OpenStudy (loser66):

now the green one is f(x) +3 https://www.desmos.com/calculator/sf239wb0wk

OpenStudy (anonymous):

did you see my graph?

OpenStudy (anonymous):

@Loser66 and what is the transformation for f( x+ 3 and f(x-2)

OpenStudy (loser66):

it says, at the beginning, the profit is higher than all 3 units. but at x =0, its profit meets the orange one and their value is 20, after then, orange one grows faster

OpenStudy (loser66):

x-2, right? we have x+2 above right, exactly the same, but x-2

OpenStudy (loser66):

that is \(f(x)=23(1.06)^{x-2}=23(1.06)^x(1.06)^{-2}=\dfrac {23(1.06)^x}{1.06)^2}=20.467(1.06)^x\)

OpenStudy (loser66):

oh, yeah, the same, just combine all in one. hehehe and you don't know how to "read" the graph, ok, it is all above.

OpenStudy (anonymous):

what is the effect though @loser66